OpenStudy (anonymous):

17 minutes after a man leaves for his morning jog, his wife leaves for a bike ride along the same road. The man jogs at an average rate of 3 mph. At what minimum rate should the wife bike to catch up with her husband within 30 minutes?

OpenStudy (lastdaywork):

Do you know the concept of relative motion ??

OpenStudy (anonymous):

i'm just not sure how to set this up into an equation.

OpenStudy (lastdaywork):

Fist tell me, are you aware of relative motion ?

OpenStudy (anonymous):

no

OpenStudy (lastdaywork):

Okay.. "At what minimum rate should the wife bike to catch up with her husband within 30 minutes?" I am not sure if the "within 30 min" is from the time when the wife leaves the home or the man. Can you please clarify ??

OpenStudy (anonymous):

30 minutes from when the wife leaves.

OpenStudy (lastdaywork):

Now, can you find the distance traveled by the man in time (17 + 3) min ??

OpenStudy (lastdaywork):

* (17 + 30) min

OpenStudy (dumbcow):

distance = rate * time man jogs for 47 min D = 3*(47/60) wife bikes for 30 min and goes same distance D = (30/60)v = 3*(47/60) solve for v

OpenStudy (anonymous):

v = 2.35

OpenStudy (dumbcow):

no (1/2) v = 2.35

OpenStudy (dumbcow):

the wife can't go slower then she will never catch up

OpenStudy (anonymous):

so it is 4.7

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yep yw

OpenStudy (triciaal):

i didn't get that distance in miles= velocity in mph times time in hours d = vt for the man d = 3(30/60) = 1.5 miles for the woman, meeting in 30 minutes she waited 17 minutes so she only has (30-17) 13 minutes to travel 1.5 miles 1.5 = v (13/60) v = 1.5(60)/13 v = 7

OpenStudy (lastdaywork):

@triciaal 30 minute period is taken from the time the wife leaves the home.

OpenStudy (triciaal):

i understand what you did but the question said she wanted to catch up with him within 30 mins so she only had the 13 minutes to bike

OpenStudy (lastdaywork):

I asked willowtrain to clarify this point; and he replied "30 minutes from when the wife leaves. "

OpenStudy (triciaal):

spanish is so much easier than english!

OpenStudy (lastdaywork):

LOL

OpenStudy (triciaal):

English is my only language, sorry